# Coreset Strikes Back: Improved Parameterized Approximation Schemes for (Constrained) $k$-Median/Means

### v2

@@ 1.1 Our Contributions
In this work, we answer 1 in the affirmative by designing a unified framework that yields EPASes for a large class of constrained $(k,z)$-clustering in metric spaces with bounded algorithmic scatter dimension. Somewhat informally, we require that the constrained $(k,z)$-clustering problem satisfies the following requirements.
@@
    There exists an algorithm $\mathcal{A}_{C}$ that computes a coreset $(Y,w)$ for the problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$.  [^5] We emphasize here that, the required bound of $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ on the coreset size is attainable even in general metric spaces for many natural constrained $(k,z)$-clustering problems; thus we do not need to rely on metric-specific properties to obtain coresets whose size independent of $\log n$.
@@
    There exists an algorithm $\mathcal{A}_{A}$ that takes as an input any weighted set of points $Z$ and candidate set of $k$ centers, $C$. Either it finds a near-optimal assignment out of all feasible (i.e., respecting all problem-specific constraints) assignments from $Z$ to $C$, or outputs that no feasible assignment from $Z$ to $C$ exists.
@@
    There exists an algorithm $\mathcal{A}_{B}$ that takes as an input a finite set of requests $Q=\{(p,r):p\in Y,r\in\mathbb{R}\}$, and an $\varepsilon\in(0,1)$, and either returns a center $c\in F$ such that (i) ${\sf d}(p,c)\leq(1+\varepsilon)r$ for all requests $(p,r)\in Q$, and (ii) $c$ satisfies problem-specific feasibility requirements; or correctly outputs that no such center exists.
We call such problem as Well Constrained $(k,z)$-Clustering problem. We allow each of the three algorithms to run in time FPT in $k,\varepsilon$, and potentially some problem-specific parameters. As we shall see later, these requirements are extremely mild and are satisfied by many natural clustering problems. Our main result is stated in the following theorem.
@@
There exists a $(1+\varepsilon)$-approximation for Well Constrained $(k,z)$-Clustering running in time $f(k,\varepsilon)\cdot|\mathcal{I}|^{O(1)}$ in metric spaces of bounded algorithmic scatter dimension, where $|\mathcal{I}|$ denotes the size of the input instance $\mathcal{I}$. In addition, if the coresets for the problem can be computed deterministically, then the algorithm is deterministic.
Now, we briefly discuss some of the consequences of our framework. A summary of these results is given in Table 1.
Constraint FPT Factor Property Theorem Section
Capacities uniform/non-uniform $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))})$ Deterministic* Thm. 3 4.1
@@
$(\alpha,\beta)$-Fairness Also: chromatic & $\ell$-diversity $\exp{(O(\frac{k\Gamma\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Gamma\Lambda}{\varepsilon}))})$ Deterministic* Thm. 4 4.2
Vanilla $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k}{\varepsilon}))})$ Randomized Thm. 5 4.3
@@
Table 1: Consequences of our EPAS framework for Well Constrained $(k,z)$-Clustering. In the second column, only the FPT part of the running time is specified, and $\Lambda\coloneqq\lambda(\varepsilon/c)$, where $\lambda(\cdot)$ is the (algorithmic) scatter dimension of the metric space, and $c$ is a constant only depends on $z$. In the second row, $\Gamma$ represents the number of equivalence classes w.r.t.  colors (see Definition 15). For the results marked as “Deterministic*”, if the coresets for the problem can be computed deterministically, then our EPAS is completely deterministic. Note that the results of [ABB+23] only work for the vanilla version, and the FPT factor of their randomized EPAS is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))$; thus our EPAS shaves off a $\log k$ factor from the exponent.
EPASes for Capacitated and Fair $(k,z)$-Clustering:
As a consequence of our framework, we obtain the first EPASes for Capacitated $(k,z)$-Clustering in metric spaces of bounded algorithmic scatter dimension. Prior to this, EPASes for Capacitated $k$-Median/Means were known only in Euclidean Spaces [CL19b, BJK18]. Furthermore, our results extend to the fair version of the problems (namely, $(\alpha,\beta)$-fair clustering [^6] In this version, points are divided into multiple (potentially overlapping) groups, and we are additionally given upper and lower bounds on the number of points of each group that can belong to each of the $k$ clusters. The goal is to find a minimum-cost clustering of points satisfying these fairness constraints. A formal definition is given in Section 4.2.  [BFS24, BGJ21]) In what follows, we state our theorems, and the key ingredients and proofs are given in Section 4.1 and Section 4.2 respectively.
@@
For any fixed $z\geq 1$, Capacitated $(k,z)$-Clustering admits an EPAS in metric spaces of algorithmic $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$ running in time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$, where $\Lambda\coloneqq\lambda({\varepsilon}/{c})$, and $c$ is a constant that only depends on $z$.
@@
Since Vanilla (a.k.a. uncapacitated/Voronoi) $(k,z)$-Clustering is a special case of Capacitated $(k,z)$-Clustering, Theorem 3 implies an EPAS for this variant with running time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$. We note that the running time of the EPAS from [ABB+23] is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))\cdot|\mathcal{I}|^{O(1)}$. Thus, note that our running time is already better than that of [ABB+23] by a factor of $\log k$ in the exponent, but we incur an additional $\log\Lambda$ factor. However, in Section 4.3, we describe a simple modification to our framework, where we replace a key deterministic branching step by a randomized one, leading to an improvement in the running time. Specifically, compared to the EPAS of [ABB+23], our resulting running time matches the dependence on $\varepsilon$, and shaves off a $\log k$ factor from the exponent. More formally, we prove the following theorem in Section 4.3.
@@
EPASes for bounded highway dimension.
We mention the following interesting corollary of our framework.
Corollary 6.
For a fixed $z\geq 1$, Well Constrained $(k,z)$-Clustering admits an EPAS in metrics of bounded highway dimension.
We find this corollary interesting due to the following reason. As mentioned earlier, metrics of bounded highway dimension are contained in the class of metrics of bounded scatter dimension. Therefore, our framework implies EPASes for Well Constrained $(k,z)$-Clustering for metric spaces of bounded highway dimension. Specifically, this includes Capacitated $(k,z)$-Clustering, which is in stark contrast with Capacitated $k$-Center, for which the result of [FV22] rules out such an EPAS (or even a PAS, i.e., an algorithm running in time $f(k)\cdot|\mathcal{I}|^{g(\varepsilon)}$).
@@ 1.2 Technical Overview
We briefly describe the technical ideas used to obtain the result, focusing mainly on constrained $k$-Median, for simplicity. Firstly, we compute a coreset $(Y,w)$ for the problem using algorithm $\mathcal{A}_{C}$, and focus our attention on obtaining a near-optimal solution on the coreset (which, in turn, implies a near-optimal solution for the original point-set). Let $\mathcal{C}=\{\mathcal{C}_{1},\dots\mathcal{C}_{k}\}$ be an optimal clustering of the coreset along with the corresponding centers $O=\{o_{1},\dots,o_{k}\}$, and let ${\sf OPT}$ be the optimal cost. We assume as have a good approximation of ${\sf OPT}$, which can be guessed approximately, up to factor $(1+\varepsilon)$, in polynomial time. On a high level, our framework maintains a a collection of $k$ request sets $(Q_{1},\dots,Q_{k})$ and a center $X=(x_{1},\dots,x_{k})$ satisfying the request sets. More specifically, each request set $Q_{i}$ is a collection of pairs of points and radii, $(p,\alpha)\in Y\times\mathbb{R}$, representing the demand that $p$ wants a center within distance $\alpha$ from itself. Additionally, the framework maintains an invariant that for $(p,\alpha)\in Q_{i}$, it must be that $p\in\mathcal{C}_{i}$, i.e., $p$ belongs to $i^{th}$ cluster in the optimal clustering $\mathcal{C}$. The main idea is to add requests to these request sets to force the algorithm refine its solution $X$ until it becomes a $(1+\varepsilon)$-approximate solution. In more detail, our framework can be conceptually broken down into multiple phases. The first phase is the leader guessing phase, which is the usual starting point in obtaining FPT-approximations for many clustering problems [CGK+19, CL19a]. In this phase, for each optimal center $o_{i}\in O$, we guess a point $p_{i}\in\mathcal{C}_{i}$ and its approximate distance [^7] Approximate up to a factor of $(1+\varepsilon)$. ${\sf d}(p_{i},o_{i})$, such that $p_{i}$ is closest to $o_{i}$ (ties broken arbitrarily). Since the size of the coreset is bounded, this guessing amounts to FPT many branches. Furthermore, this gives a baseline of minimum distances in each of the $k$ clusters in $\mathcal{C}$ that will be used by the algorithm later. Next, we add requests $(p_{i},d(p_{i},o_{i}))$ to $Q_{i}$ for all $i\in[k]$, and using algorithm $\mathcal{A}_{B}$, we compute a feasible set of centers $X$ satisfying all $Q_{i}$s. This means that, for each $i\in[k]$, center $x_{i}\in X$ satisfies ${\sf d}(p_{i},x_{i})\leq(1+\varepsilon){\sf d}(p,o_{i})$. Finally, we use algorithm $\mathcal{A}_{A}$ to compute a near-optimal assignment of points in $Y$ to $X$, to obtain a corresponding clustering $\mathcal{C}^{\prime}=(\mathcal{C}^{\prime}_{1},\dots,\mathcal{C}^{\prime}_{k})$. Then, as long as the cost of the current clustering $\mathcal{C}^{\prime}$ is larger than $(1+\varepsilon)\cdot{\sf OPT}$, we argue that there must exist an unhappy pair $(p,i)$, where $p\in\mathcal{C}_{i}$, i.e., $p$ is assigned to $o_{i}$ in the optimal solution, such that ${\sf d}(p,x_{i})>(1+\varepsilon/3){\sf d}(p,o_{i})$. Since, the size of coreset $Y$ is small, we can guess such unhappy pair $(p,i)$ using at most $|Y|\cdot k$ many branches. Next, we add the request $(p,\frac{{\sf d}(p,x_{i})}{1+\varepsilon/3})$ to request set $Q_{i}$. Finally, we recompute $X$ using $\mathcal{A}_{B}$ to additionally satisfy the new request, and repeat the whole process, if the cost of the recomputed $X$ is still larger than $(1+\varepsilon){\sf OPT}$. By carefully analyzing the nature of such unhappy pairs, we show the crucial property that for each request set $Q_{i}$, the radii of requests in $Q_{i}$ lie in an interval of width bounded by $O(1/\varepsilon)$. This, in turn, allows us to bound the number of requests that can be added to each $Q_{i}$ in metric spaces with bounded algorithmic scatter dimension. Thus, this leads to a recursive algorithm with branching factor $|Y|\cdot k$ and depth $O(\frac{k}{\varepsilon}\log(\frac{1}{\varepsilon})\cdot\lambda(\frac{\varepsilon}{c}))$, where $\lambda(\cdot)$ is the algorithmic scatter dimension of the metric space, and $c$ is a constant. Since $|Y|=\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$, this leads to the claimed running time by employing the standard trick, namely, $(\log n)^{t}=t^{O(t)}\cdot n^{O(1)}$.
Comparison to existing frameworks.
On a high level, our algorithm follows the outline of the framework recently proposed by [ABB+23], that unifies existing EPASes for Voronoi clustering problems. However, we face multiple challenges when we try to extend their framework to our problems. For instance, [ABB+23] introduce a notion of upper bounds, and devise them for their problems, that help bound the aspect ratio of radii in $Q_{i}$s. However, these upper bounds are inherently based on Voronoi properties, and it is not clear how to design them for our problems. This is where, we exploit the existence of general metric coresets. As mentioned before, coresets allows us to get a handle on the leaders and their distances, which in turn, allows us to bound the aspect ratio of radii in each request set, the task that upper bounds achieved in [ABB+23]. Another bottleneck that is apparent when extending their framework to our problems, is finding unhappy pairs. As mentioned before, if the cost of current solution $(X,\mathcal{C}^{\prime})$ is larger than $(1+\varepsilon){\sf OPT}$, there exists a unhappy pair, and we again exploit the compact size of coresets to find one. This is not a bottleneck for Voronoi clustering problems, since it can be shown that, in this case, there is a unhappy point $p\in Y$ such that ${\sf d}(p,X)>(1+\varepsilon/3){\sf d}(p,O)$. [^8] For a given center set $S$ and a point $p\in Y$, ${\sf d}(p,S):=\min_{s\in S}{\sf d}(p,s)$. This means that, if $p$ was assigned to $o_{i}$ in the optimal solution, then ${\sf d}(p,x_{i})\geq d(p,X)>(1+\varepsilon/3){\sf d}(p,o_{i})$, and hence $(p,i)$ is an unhappy pair, as required. Therefore, we can add a new request $(p,\tfrac{{\sf d}(p,x_{i})}{1+\varepsilon})$ to $Q_{i}$, so that the algorithm refines $x_{i}$ to satisfy the new request. However, for constrained clustering, it is possible that $p$ is assigned to $x_{j}\in X$ in $\mathcal{C}^{\prime}$, while it is assigned to $o_{i}$ in $\mathcal{C}$, for $i\neq j$, and it holds that ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$ due to the cost of $(X,\mathcal{C}^{\prime})$. Furthermore, it can happen that ${\sf d}(p,x_{i})$ is much smaller than ${\sf d}(p,x_{j})$ (the worst case is when ${\sf d}(p,x_{i})=0$, while ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$), which means that adding a request $(p,\tfrac{{\sf d}(p,x_{i})}{1+\varepsilon})$ does not refine $x_{i}$, as $x_{i}$ might have already satisfied the request. This can potentially make the algorithm run forever.
As mentioned earlier, coresets enable us to initialize the request sets $Q_{i}$ much more accurately due to the leader guessing phase. This, along with a crucial observation on the nature of unhappy pairs, allows us to bound the radii aspect ratio of requests in each $Q_{i}$ by $O(1/\varepsilon)$, whereas the loose upper bound approach of [ABB+23] bounds the radii aspect ratio only by $O(k/\varepsilon)$. This helps us in obtaining improved running time for uncapacitated $(k,z)$-clustering, as described above.

### v3

@@ 1.3 Our Contributions
In this subsection, we begin by presenting an overview of our unified framework for constrained clustering problems, followed by a statement of the main result associated with this framework. We then discuss a range of applications of this result.
@@
In this work, we design a unified framework that yields EPASes for a large class of constrained $(k,z)$-clustering problems, in metric spaces with bounded (algorithmic) scatter dimension. This, in particular, answers 1 of Abbasi et al. [ABB+23] from FOCS 2023. Towards this, we say that a center set $X$ of size $k$ is feasible if it satisfies the center constraints Analogously, we say, for a center set $X$ and an assignment $f$, $(X,f)$ is feasible if $X$ is feasible and the assignment $f$ from from $P$ to $X$ satisfies the assignment constraint. We require that the constrained $(k,z)$-clustering problem satisfies the following (somewhat informally stated) requirements. See Section 3.1 for the formal setup.
@@
    There exists an algorithm $\mathcal{A}_{C}$ that computes a (client set) coreset $(Y,w)$ for the problem of size $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$.  [^8] We emphasize here that, the required bound of $\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$ on the coreset size is attainable even in general metric spaces for many natural constrained $(k,z)$-clustering problems; thus we do not need to rely on metric-specific properties to obtain coresets whose size is independent of $\log n$.
@@
    There exists an algorithm $\mathcal{A}_{A}$ that takes as an input any weighted set of points $Z$ and candidate set of $k$ centers, $C$. Either it finds a near-optimal assignment from all feasible assignments from $Z$ to $C$, or outputs that no feasible assignment from $Z$ to $C$ exists.
@@
[…]
We call such a problem as Well Constrained $(k,z)$-Clustering problem. We allow each of the three algorithms to run in time FPT in $k,\varepsilon$, and potentially some problem-specific parameters. These requirements are extremely mild and are satisfied by many natural clustering problems. For instance, the capacitated $(k,z)$-clustering problem admits a coreset of the aforementioned size [CL19] in general metrics, its assignment algorithm can be implemented via a minimum-cost flow subroutine, and the algorithm $\mathcal{A}_{B}$ returns a feasible center with maximum capacity.
Our main result is summarized in the following theorem.
@@
There exists a $(1+\varepsilon)$-approximation for Well Constrained $(k,z)$-Clustering running in time $h(k,\varepsilon)\cdot|\mathcal{I}|^{O(1)}$ in metric spaces of bounded scatter dimension, where $|\mathcal{I}|$ denotes the size of the input instance $\mathcal{I}$. In addition, if the coresets for the problem can be computed deterministically, then the algorithm is deterministic.
As a corollary of this theorem, we obtain EPASes for several well-constrained $(k,z)$-clustering problems in metric spaces of bounded scatter dimension. These results are summarized in the first half of Table 1. In Section 1.3.1, we state the results and outline the key ingredients used to obtain them. The formal proofs are given in Section 4.1.1, Section 4.2.1, and Section 4.1.3, respectively.
In light of 1, it is natural to ask whether Theorem 1 can be extended to metric spaces of algorithmic scatter dimension. While there are fundamental obstacles to achieving such an extension in full generality, which we discuss in Section 1.4, we are nevertheless able to obtain EPASes for certain specific instances of Well Constrained $(k,z)$-Clustering. These results are stated in the second half of Table 1, and in Section 1.3.2, with a detailed discussion provided in Section 4.2.
Dimension Constraint FPT Factor Theorem Section
Scatter Dimension Capacities uniform/non-uniform $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))})$ * Thm. 2 4.1.1
Matroid 4.1.2
Fault Tolerant 4.1.3
@@
Algorithmic Scatter Dimension $(\alpha,\beta)$-Fairness Also: chromatic & $\ell$-diversity $\exp{(O(\frac{k\Gamma\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Gamma\Lambda}{\varepsilon}))})$ * Thm. 3 4.2.1
Vanilla $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k}{\varepsilon}))})$ Thm. 4 4.2.2
@@
Table 1: Consequences of our EPAS framework for Well Constrained $(k,z)$-Clustering. In the third column, only the FPT part of the running time is specified, and $\Lambda\coloneqq\lambda(\varepsilon/c)$, where $\lambda(\cdot)$ is the (algorithmic) scatter dimension of the metric space, and $c$ is a constant that only depends on $z$. In the fourth row, $\Gamma$ represents the number of equivalence classes w.r.t.  colors (see Definition 17). For the results marked as “*”, if the coresets for the problem can be computed deterministically, then our EPAS is completely deterministic. Note that the results of [ABB+23] only work for the vanilla version, and the FPT factor of their randomized EPAS is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))$; thus our EPAS shaves off a $\log k$ factor from the exponent.
1.3.1 EPASes for Bounded Scatter Dimension
Although, the problems mentioned here implicitly impose a soft constraint of selecting distinct centers in the solution, this can be circumvented by the method of color coding (see Section 4.1). This technique preprocesses the input instance in FPT time such that the facilities have one of the $k$ colors and for a fixed optimal center set, all the centers have distinct colors.
EPASes for Capacitated, Matroid, and Fault Tolerant $(k,z)$-Clustering:
As a consequence of our framework, we obtain the first EPASes for Capacitated $(k,z)$-Clustering in metric spaces of bounded scatter dimension. Prior to this, EPASes for Capacitated $k$-Median/Means were known only in Euclidean Spaces [CL19, BJK18]. Furthermore, our results extend to Matroid and Fault tolerant $(k,z)$-Clustering.
@@ 1.3.1 EPASes for Bounded Scatter Dimension
For any fixed $z\geq 1$, the following problems admit an EPAS in metric spaces of $\varepsilon$-scatter dimension bounded by $\lambda(\varepsilon)$ running in time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$, where $\Lambda\coloneqq\lambda({\varepsilon}/{c})$, and $c$ is a constant that only depends on $z$: (i) Capacitated $(k,z)$-Clustering, (ii) Matroid $(k,z)$-Clustering, and (iii) Fault tolerant $(k,z)$-Clustering.
EPASes for metric spaces of bounded highway dimension.
As referenced earlier, it has been observed that metric spaces of bounded highway dimension (BHD) also have bounded scatter dimension. Assuming this result, our framework implies an EPAS for Well Constrained $(k,z)$-Clustering for any fixed $z\geq 1$ in BHD metric spaces; although we refrain from stating the result as a formal corollary since a formal proof of the claim that BHD metric spaces have bounded scatter dimension is not available in the public domain.
We find this result interesting for a couple of reasons. Firstly, we are not aware of coresets of size $O_{\varepsilon,k}(1)$ for Well Constrained $(k,z)$-Clustering in BHD metric spaces, i.e., coresets whose size is independent of $n$. Therefore, this would be the first EPAS for many instances of Well Constrained $(k,z)$-Clustering in BHD metric spaces, to the best of our knowledge. Secondly, our result in particular implies an EPAS for Capacitated $(k,z)$-Clustering in BHD metrics. This result is in stark contrast with Capacitated $k$-Center, for which the result of [FV22] rules out such an EPAS (or even a PAS, i.e., an algorithm running in time $f(k)\cdot|\mathcal{I}|^{g(\varepsilon)}$) in BHD metrics.
1.3.2 EPASes for Bounded Algorithmic Scatter Dimension
We now show applications of our framework to obtain EPAS for metrics of bounded algorithmic scatter dimension for certain special cases.
Fair $(k,z)$-Clustering.
Our framework extends to handle $(\alpha,\beta)$-fair clustering [^9] In this version, points are divided into multiple (potentially overlapping) groups, and we are additionally given upper and lower bounds on the number of points of each group that can belong to each of the $k$ clusters. The goal is to find a minimum-cost clustering of points satisfying these fairness constraints. A formal definition is given in Section 4.2.1. [BFS24a, BJK18]). Again, this problem has a soft constraint of having distinct centers in $X$, but note that if the clusters corresponding to multiple copies of a center are ”fair”, then they can be merged together by keeping just one copy of the center.
@@ 1.3.2 EPASes for Bounded Algorithmic Scatter Dimension
Since Vanilla (a.k.a. uncapacitated/Voronoi) $(k,z)$-Clustering is a special case of Well constrained $(k,z)$-Clustering, we obtain an EPAS for this variant with running time $\exp{(O(\frac{k\Lambda}{\varepsilon}\log(\frac{1}{\varepsilon})\log(\frac{k\Lambda}{\varepsilon}))}\cdot|\mathcal{I}|^{O(1)}$. We note that the running time of the EPAS from [ABB+23] is $\exp(O(\frac{k\Lambda}{\varepsilon}\log^{2}(\frac{k}{\varepsilon})))\cdot|\mathcal{I}|^{O(1)}$. Thus, note that our running time is already better than that of [ABB+23] by a factor of $\log k$ in the exponent, but we incur an additional $\log\Lambda$ factor. However, in Section 4.2.2, we describe a simple modification to our framework, where we replace a key deterministic branching step by a randomized one, leading to an improvement in the running time. Specifically, compared to the EPAS of [ABB+23], our resulting running time matches the dependence on $\varepsilon$, and shaves off a $\log k$ factor from the exponent. More formally, we prove the following theorem in Section 4.2.2.
@@ 1.4 Technical Overview
We start with an overview of the main ideas behind the EPAS of Theorem 1, including the new challenges faced to handle constrained clustering, as compared to prior work, and how we handle them. This is discussed in Section 1.4.1. Then, in Section 1.4.2, we discuss why we cannot extend the result to algorithmic scatter dimension. Finally, in Section 1.4.3 we compare our results with another orthogonal approach based on the work of [HLL+25].
1.4.1 Our Techniques
@@
We briefly describe the technical ideas used to obtain the result, focusing mainly on constrained $k$-Median, for simplicity. Our framework is roughly follows the outline of the framework of [ABB+23] proposed for Voronoi clustering problems. On a high level, their framework is as follows. Recall that, in their problem, a point is assigned to the closest center, unlike our problem, where a point can be assigned to multiple centers. Let $\mathcal{C}=\{\mathcal{C}_{1},\dots\mathcal{C}_{k}\}$ be an optimal clustering of the coreset along with the corresponding centers $O=\{o_{1},\dots,o_{k}\}$, and let $\sf OPT$} be the optimal cost. The main idea of their framework is to maintain a collection of $k$ request sets $(Q_{1},\dots,Q_{k})$ and a set $X=(x_{1},\dots,x_{k})$ of centers satisfying the request sets. More specifically, each request set $Q_{i}$ is a collection of point-distance pairs $(p,\alpha)\in Y\times\mathbb{R}$, representing the demand that $p$ wants a center within distance $\alpha$ from itself. Additionally, the framework maintains an invariant that for $(p,\alpha)\in Q_{i}$, it must be that $p\in\mathcal{C}_{i}$, i.e., $p$ belongs to $i^{th}$ cluster in the optimal clustering $\mathcal{C}$. The idea is to add requests to these request sets to force the algorithm to refine its solution $X$ until it becomes a $(1+\varepsilon)$-approximate solution. To argue this refinement, they show that if the cost of the current clustering $\mathcal{C}^{\prime}$ is larger than $(1+\varepsilon)\cdot{\sf OPT}$, then there must exist an unhappy point $p$: $p\in\mathcal{C}_{i}$, i.e., $p$ is assigned to $o_{i}$ in the optimal solution, such that ${\sf d}(p,x_{i})>(1+\varepsilon/3)\cdot{\sf d}(p,o_{i})$, based on the fact that $p$ is assigned to the closest center. This gives a handle to refine $x_{i}$ by adding the request $(p,\tfrac{d(p,X)}{1+\varepsilon})$ to $Q_{i}$, and by guessing $i\in[k]$ correctly. [^10] One may view this as the algorithm trying to simulate $o_{i}$ by refining $x_{i}$. Now, if the aspect ratio of the distances in $Q_{i}$ are bounded, then it can be shown that such a refinement can occur only bounded number of times, when the metric space has bounded algorithmic scatter dimension. Towards this, they introduce a concept of upper bounds to initialize each $Q_{i}$, which again, is based on Vornoi-assignment. However, we face multiple challenges when we try to extend their framework to our problems.
  • Upper bounds: As mentioned above, the upper bounds are inherently based on Vornoi properties, and it is not clear how to design them for constrained problems.
  • Assigning a point to multiple centers: This distinction highlights a fundamental difference between our setting and the one considered in [ABB+23]. In their problem, a point $p$ is always assigned to a (closest) single center. This allows them to guess the index of its center $o_{i}$ in the optimal solution, and then use its current distance to refine $x_{i}$, as mentioned above. On the other hand, in our setting, each point $p$ may be assigned to multiple centers. In fact, the bottleneck arises in a subtle way even when a point is assigned to only one (not necessarily a closest one) center. This is because, for constrained clustering, it is possible that $p$ is assigned to $x_{j}\in X$ in $\mathcal{C}^{\prime}$, while it is assigned to $o_{i}$ in $\mathcal{C}$, for $i\neq j$, and it holds that ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$ due to the cost of $(X,\mathcal{C}^{\prime})$. On the other hand, it can happen that ${\sf d}(p,x_{i})$ is much smaller than ${\sf d}(p,x_{j})$ (the worst case is when ${\sf d}(p,x_{i})=0$, while ${\sf d}(p,x_{j})>(1+\varepsilon/3){\sf d}(p,o_{i})$), which means that adding a request $(p,\tfrac{{\sf d}(p,x_{i})}{1+\varepsilon})$ does not refine $x_{i}$, as $x_{i}$ might have already satisfied the request. This can potentially make the algorithm run forever.
We now highlight our ideas to tackle these challenges.
Bounding distance aspect ratio of requests. Towards this goal, we exploit the existence of small sized coresets for the problem: we use the leader guessing technique, which is a usual starting point in obtaining FPT-approximations for many clustering problems [CGK+19b, CL19]. Let $(Y,w)$ be the coreset, and $(O,f^{*})$ be its optimal solution. More specifically, for each optimal center $o_{i}\in O$, we guess a point $p_{i}\in P$ with $f^{*}(p,o_{i})>0$ and its approximate distance $\vartheta_{i}\approx{\sf d}(p_{i},o_{i})$ [^11] Approximate up to a factor of $(1+\varepsilon)$., such that $p_{i}$ is closest to $o_{i}$ (ties broken arbitrarily). Since the size of the coreset is bounded, this guessing amounts to FPT many branches. Next, we add requests $(p_{i},\vartheta_{i})$ to $Q_{i}$ for all $i\in[k]$, and using algorithm $\mathcal{A}_{B}$, we compute a feasible set of centers $X$ satisfying all $Q_{i}$s. This means that, for each $i\in[k]$, center $x_{i}\in X$ satisfies ${\sf d}(p_{i},x_{i})\leq(1+\varepsilon){\sf d}(p,o_{i})$. Finally, we use algorithm $\mathcal{A}_{A}$ to compute a near-optimal assignment $f$ of points in $Y$ to $X$. This gives a baseline of minimum distances in each of the $k$ request sets maintained by the algorithm, and hence, allows us to bound the distance aspect ratio of requests in each $Q_{i}$ by $O(1/\varepsilon)$, whereas the loose upper bound approach of [ABB+23] bounds it only by $O(k/\varepsilon)$. This helps us in obtaining improved running time for uncapacitated $(k,z)$-clustering, as described above.
Handling non-Voronoi assignments. To tackle the multiple assignments to (non-Voronoi) centers issue, we consider a different line of argument. Specifically, we do the following. Let $(X,f)$ be the current solution whose cost is more than $(1+\varepsilon)$ times the optimal cost. Let $(O,f^{*})$ be a fixed optimal solution. Let $\hat{f}^{*}$ be the assignment of points in $P$ to $X$ based on $f^{*}$, i.e., $\hat{f}^{*}(p,x_{i}):=f^{*}(p,o_{i})$, for all $i\in[k]$. Our crucial observation is that there must exist a pair $(p,i)$, called unhappy pair, such that $f^{*}(p,x_{i})>0$, however $d(p,x_{i})>(1+\varepsilon/3)d(p,o_{i})$; otherwise, this implies that the cost $(X,\hat{f}^{*})$ is at most $(1+\varepsilon)$ times optimal cost, and hence a contradiction to the cost of $(X,f)$, since $f$ is the best possible assignment on $X$. Since, the size of coreset $Y$ is small, we can guess such unhappy pair $(p,i)$ using at most $|Y|\cdot k$ many branches. Next, we add the request $(p,\frac{{\sf d}(p,x_{i})}{1+\varepsilon/3})$ to request set $Q_{i}$. Finally, we recompute $X$ using $\mathcal{A}_{B}$ to additionally satisfy the new request, and repeat the whole process, if the cost of the recomputed $X$ is still larger than $(1+\varepsilon){\sf OPT}$. From the algorithmic point of view, the key step is to always maintain a set of centers that not only satisfies all the request sets, but also satisfies the additional constraints of the problem. For example, in the case of capacitated clustering, in every iteration, we maintain a center $x_{i}$ that has the maximum capacity satisfying the request set $Q_{i}$, for every $i\in[k]$. This allows us to leverage the scatter dimension bound, while maintaining a feasible solution during the execution of the algorithm. Furthermore, we show that these ideas allow us to bound the distances aspect ratio of requests in each $Q_{i}$ by $O(1/\varepsilon)$. This, in turn, allows us to bound the number of requests that can be added to each $Q_{i}$ in metric spaces with bounded scatter dimension. Thus, this leads to a recursive algorithm with branching factor $|Y|\cdot k$ and depth $O(\frac{k}{\varepsilon}\log(\frac{1}{\varepsilon})\cdot\lambda(\frac{\varepsilon}{c}))$, where $\lambda(\cdot)$ is the algorithmic scatter dimension of the metric space, and $c$ is a constant. Since $|Y|=\left(\frac{k\log n}{\varepsilon}\right)^{O(1)}$, this leads to the claimed running time by employing the standard trick, namely, $(\log n)^{t}=t^{O(t)}\cdot n^{O(1)}$.
1.4.2 Challenges of Extending to Algorithmic Scatter Dimension
Having presented the technical overview of our EPAS for metrics of bounded scatter dimension, we now revisit the issue of why this framework cannot be extended to metrics of bounded algorithmic scatter dimension. The notion of algorithmic scatter dimension is tied to a specific ball-intersection algorithm that returns a center satisfying a given set of distance constraints. Its boundedness guarantee applies only to this particular procedure; alternative algorithms may result in unbounded behavior.
In constrained clustering, however, center selection is subject to additional feasibility requirements. For example, capacitated clustering enforces global capacity constraints, while matroid constraints restrict the center set to be an independent set. The ball-intersection algorithm underlying algorithmic scatter dimension is oblivious to such constraints, and its output may violate feasibility. Furthermore, many constraints impose exclusivity or soft restrictions on center selection, such as selecting each facility at most once. Modifying the ball-intersection algorithm to respect these constraints may invalidate the boundedness guarantees of algorithmic scatter dimension.
These issues are intrinsic to constrained clustering. To overcome them, we base our framework on the purely combinatorial notion of $\varepsilon$-scatter dimension. This relaxation allows us to reason about arbitrary feasible center constructions while still exploiting bounded scattering behavior.
1.4.3 Comparison to the work of Huang et al. [HLL+25].